give the coordinates of the vertex
y=x^2+4x+1

Answer by nerdybill(7384)

give the coordinates of the vertex
y=x^2+4x+1
.
coordinate of the x-axis:
x = -b/(2a)
x = -4/(2(1))
x = -4/2
x = -2
.
to find coordinate of the y-axis, plug it back in to the original equation:
y=x^2+4x+1
y=(-2)^2+4(-2)+1
y=4+4(-2)+1
y=4+(-8)+1
y=4-8+1
y=-4+1
y=-3
.
Vertex is at (-2,-3)